Find the common factors of the given terms.
 3x^2y^3 , 10x^3y^2 , 6x^2 y^2z

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 3x^2y^3, 10x^3y^2and 6x^2y^2z
3x^2y^3= 3 × x × x × y × y × y
10x^3y^2= 2 × 5 × x × x × x × y ×y
6x^2y^2z = 3 × 2 × x × x × y × y × z
Common factors of 3x^2y^3, 10x^3y^2and 6x^2y^2z are x^2, y^2and, x^2×y^2= x^2y^2

Thanks